OpenStudy (anonymous):

A golf ball is struck with a club with a force of 1500N. The time of contact between the golf ball and the club is 20ms. The mass of the ball is 0.2kg How high will the golf ball will rise?

OpenStudy (anonymous):

@welshfella @hartnn any idea?

OpenStudy (anonymous):

I've calculated the initial velocity using the Impulse formula so I = F*t I = m(v-u) F*t = m(v-u) 1500*(20*10^-3) = 0.2(v-0) 30 = 0.2v 150 = v so I know the initial velocity now, but I don't quite understand exactly how am I supposed to calculate the max height from the projectile of the golf ball